The following pH data were collected in the analysis of water from an upstate New York lake as part of an acid rain study: 4.17, 4.20, 4.19, 4.23, 4.22, 4.14, 4.20, and 4.15. Calculate the following.



(a) mean


(b) median


(c) percent average deviation


(d) percent standard deviation

1
Expert's answer
2022-04-26T05:41:04-0400

(a) Mean value = (4.17 + 4.20 + 4.19 + 4.23 + 4.22 + 4.14 + 4.20 + 4.15) / 8 = 4.1875

(b) Median value = (4.23 + 4.22) / 2 = 4.225 (even number of data)

(c) Percent average deviation = (|4.17 - 4.1875| + |4.20 - 4.1875| + |4.19 - 4.1875| + |4.23 - 4.1875| + |4.22 - 4.1875| + |4.14 - 4.1875| + |4.20 - 4.1875| + |4.15 - 4.1875|) / 8 × 100% = 0.205 / 8 × 100% = 2.56%

(d) Percent standard deviation = √{(4.17 - 4.1875)2 + (4.20 - 4.1875)2 + (4.19 - 4.1875)2 + (4.23 - 4.1875)2 + (4.22 - 4.1875)2 + (4.14 - 4.1875)2 + (4.20 - 4.1875)2 + (4.15 - 4.1875)2} / 8 = 0.030 or 3.0%